# Unsupervised Representation Learning from Sparse Transformation Analysis

### v2

@@ 4.3 Fluid-Dynamic OT for Posterior Flow
Theorem 4.1 (Benamou-Brenier Formula (BBF) [benamou2000computational]).
@@ 4.4 Brownian Motion for Latent Prior Evolution
In line with [song2023flow], as we do not assume any prior knowledge of each transformation, we would like to enforce minimally informative priors. This can be achieved by considering the time evolution as Brownian motion, i.e., random trajectories. To this end, we define the potential function $\psi^{k}({\bm{z}})=-D_{k}\log p({\bm{z}}_{t})$ which advects the density $p({\bm{z}})$ through the induced velocity field $\nabla\psi^{k}({\bm{z}})$. Then according to the continuity equation, the prior evolves as:
@@ 4.5 Latent Posterior Time Evolution
TABLE II: Equivariance error $\mathcal{E}_{k}$ and average log-likelihood $\log p({\bm{x}}_{t})$ on Shapes3D [3dshapes18].
@@
Figure 5: Traversals using each individual learned flow field on Shapes3D [3dshapes18]. In the bracket, we indicate the transformation which the traversal results look most like. Each latent flow has separate samples per row transforming from left to right.
@@ 5.1.1 Datasets
We evaluate our method on two widely-used benchmarks for standard representation learning, namely MNIST [lecun1998mnist] and Shapes3D [3dshapes18]. For MNIST, The basic ‘pure’ transformations consist of Scaling, Rotation, and Coloring. For Shapes3D, we use the self-contained four transformation primitives, including Floor Hue, Wall Hue, Object Hue, and Scale. On both datasets, we use our spike and slab prior to generate sequences that are composed of ‘composite’ transformations.
Beyond the toy datasets, we also evaluate our method on challenging Falcol3D and Issac3D [nie2020semi], two complex large-scale and real-world datasets that contain sequences of different transformations. Specifically, Falcol3D consists of indoor 3D scenes with different lighting conditions and camera positions, while Isaac3D is comprised of various robot-arm movements in dynamic environments. Since the image sequences are short, we do not consider speed variations but only enforce the spike prior to generate data sequences with sparsely-varying transformations.
We further conduct some preliminary experiments of applying our method to real-world video analysis, including autonomous driving videos on Cityscape [cordts2016cityscapes] and behavior videos of social agents on CalMS [sun2021multi]. Different from the used datasets above, we directly feed raw video sequences as input and let the model discover independent motions.
@@ 5.1.2 Baselines
We compare our method with some representative approaches in the field of disentangled and equivariant representation learning, including LatentFlow [song2023flow] and PoFlow [song2023latent] which adopt potential flow to evolve the latent samples, Topographic VAE (TVAE) [keller2021topographic] which posses topographic structured latent space, SlowVAE [klindt2021towards] which proposes the sparse Laplacian prior $p({\bm{z}}_{t}|{\bm{z}}_{t-1})=\prod\nicefrac{{\alpha\lambda}}{{2\Gamma(1/\alpha)}}\exp{(-\lambda|z_{t,i}-z_{t-1,i}|^{\alpha})}$, and $\beta$-VAE [higgins2016beta] and FactorVAE [kim2018disentangling] which encourage the factorization of the single dimensions of latent samples. We also use the vanilla VAE [kingma2013auto] as a controlled baseline.
@@ 5.1.3 Metrics
We mainly evaluate the baselines using the equivariance error which is defined as $\mathcal{E}_{k}{=}\sum_{t=1}^{T}|{\bm{x}}_{t}-\texttt{Decode}({\bm{z}}_{t})|$ where ${\bm{x}}_{t}$ is the element of sequences of each transformation primitive (e.g., scaling and rotation). Since our method is unsupervised, we inspect the traversal results of each basic vector field $\nabla u^{k}+{\bm{r}}^{k}$ and select the index whose flow looks the most like the target transformation. The average log-likelihood of the sequence is also evaluated on the test set. Besides these two metrics, we also adopt the metric Variational Predictability (VP) score [zhu2020learning] to evaluate the disentanglement performance. Readers are kindly referred to Sec. D.3 of the supplementary for these results.
@@ 5.2.2 Quantitative Results
Table I and II present the evaluation results of the equivariance error and log-likelihood on MNIST and Shapes3D, respectively. We see STA achieves very competitive performance against other baselines. Specifically, STA outperforms all the unsupervised approaches by a large margin on equivariance error and rivals PoFlow [song2023latent] which requires supervision of each transformation primitive. Moreover, our method yields the highest log-likelihood on the test set, which is likely accounted for by the fact that our method incorporates a sophisticated transformation-centric prior over latent states which matches the statistics of the data. The sparse combination of multiple transformations thus can be seen as a kind of data augmentation. Among the transformations, the rotation has the smallest equivariance error. We expect this gain is largely due to the rotational vector field ${\bm{r}}({\bm{z}})$ introduced by the Helmholtz decomposition.
@@ 5.3.3 Periodic Transformations
Figure 10: Traversals using each individual learned flow field on Falcol3D [nie2020semi]. In the bracket, we indicate the transformation which the traversal results look most like. Each latent flow has separate samples per row transforming from left to right. The bottom row displays the traversal result generated by the $6$’th latent flow field.
@@
TVAE [keller2021topographic] 11477.81 12568.32 11807.34 11829.33 11539.69 11736.78 11951.45
PoFlow [song2023latent] 8312.97 7956.18 8519.39 8871.62 8116.82 8534.91 8994.63
LatentFlow [song2023flow] 5798.42 6145.09 6334.87 6782.84 6312.95 6513.68 6614.27
@@ 5.5 Real-world Video Analysis
Figure 11: Traversals using each individual learned flow field on Issac3D [nie2020semi]. In the bracket, we indicate the transformation which the traversal results look most like. Each latent flow has separate samples per row transforming from left to right.
@@
TVAE [keller2021topographic] 8441.65 8348.23 8495.31 8251.34 8291.70 8741.07 8456.78 8512.09
PoFlow [song2023latent] 6572.19 6489.35 6319.82 6188.59 6517.40 6712.06 7056.98 6343.76
LatentFlow [song2023flow] 3659.72 3993.33 4170.27 4359.78 4225.34 4019.84 5514.97 3876.01
@@ 5.5.1 Robot Arms and Indoor Scenes
Fig. 10 and 11 show the learned latent flows on Falcol3D and Issac3D [nie2020semi], respectively. As can be seen from the figures, even on these challenging large-scale datasets, our method still allows for unsupervised disentanglement of complex real-world transformations. Table V and VII compares the equivariance error on the two datasets. Similar to the results on MNIST and Shapes3D, our method achieves very competitive performance against supervised ones. This demonstrates that the proposed sparsity priors also scale up to sequences of complex transformations.
Figure 12: Traversal results of learned latent flows on CalMS [sun2021multi]. For each latent flow, we display two exemplary sequences, and the flow transforms the image from left to right.
@@ 5.5.2 Agent Behavioral Videos
TABLE VIII: Behavior classification results on CalMS [sun2021multi].
Method MARS [segalin2021mouse] B-Kind [sun2021multi] Trajectory-LSTM [sun2021multi] STA
@@
Figure 14: Traversals results of learned flow fields on downsampled segmentation masks of Cityscape [cordts2016cityscapes]. Each latent flow transforms the image from left to right.
We apply our method to disentangle the complex social interactions of mice on Task $1$ of CalMS [sun2021multi]. On this dataset, there exist three ground truth interactions, namely ‘investigation’, ‘attack’, and ‘mount’. We thus define $3$ latent flows and let the model learn the interactions from the raw videos. The images are of the resolution $128\times 128$.
@@ 5.5.3 Autonomous Driving Videos
Finally, we take a step further to evaluate our method on Cityscape [cordts2016cityscapes], the challenging real-world autonomous driving videos. We take the sequences of segmentation masks as the training data and downsample the resolution to $64{\times}64$. Fig. 14 displays some exemplary traversals of different latent flows. On this dataset, there are no ground truth generative factors so we may have some reasonable interpretations according to the disentangled transformations: we may interpret latent flow 0 as turning left (the sidewalk region on the right side shrinks), latent flow $1$ as getting closer to the front car (the car region expands), latent flow $2$ as getting away from the front car (the car region shrinks and disappears), and latent flow $3$ as changing the right side from terrain to sidewalk. Notice that this is an initial attempt to apply our method to complex real-world video analysis. Nonetheless, this preliminary experiment demonstrates that our method could have real-world applicability for video understanding.
@@ 6 Conclusion
[Uncaptioned image] Yue Song received the B.Sc. cum laude from KU Leuven, Belgium and the joint M.Sc. summa cum laude from the University of Trento, Italy and KTH Royal Institute of Technology, Sweden, and the Ph.D. summa cum laude from the Multimedia and Human Understanding Group (MHUG) at the University of Trento, Italy. Currently, he is a post-doctoral research associate at Caltech. His research interests are structured representation learning.
[Uncaptioned image] T. Anderson Keller is a postdoctoral research fellow at the Kempner Institute at Harvard University. He completed his doctorate under the supervision of Max Welling at the University of Amsterdam. His current research focuses on structured representation learning, probabilistic generative modeling, and biologically plausible learning. His research explores ways to develop deep probabilistic generative models that are meaningfully structured with respect to observed, real-world transformations.
[Uncaptioned image] Yisong Yue is a professor of Computing and Mathematical Sciences at Caltech. My research interests lie primarily in machine learning, and span the entire theory-to-application spectrum from foundational advances all the way to deployment in real systems. I work closely with domain experts to understand the frontier challenges in applied machine learning, distill those challenges into mathematically precise formulations, and develop novel methods to tackle them.
[Uncaptioned image] Pietro Perona received the graduate degree in electrical engineering from the Università di Padova in 1985 and the PhD degree in electrical engineering and computer science from the University of California at Berkeley in 1990. After a postdoctoral fellowship at MIT in 1990-1991 he joined the faculty of the California Institute of Technology, Caltech in 1991, where he is now an Allen E. Puckett professor of electrical engineering and computation and neural systems. His current interests include visual recognition, modeling vision in biological systems, modeling and measuring behavior, and Visipedia. He has worked on anisotropic diffusion, multiresolution-multiorientation filtering, human texture perception and segmentation, dynamic vision, grouping, analysis of human motion, recognition of object categories, and modeling visual search.
[Uncaptioned image] Max Welling is a research chair in Machine Learning at the University of Amsterdam and a Distinguished Scientist at MSR. He is a fellow at the Canadian Institute for Advanced Research (CIFAR) and the European Lab for Learning and Intelligent Systems (ELLIS) where he also serves on the founding board. His previous appointments include VP at Qualcomm Technologies, professor at UC Irvine, postdoc at U. Toronto, and UCL under the supervision of Prof. Geoffrey Hinton, and postdoc at Caltech under the supervision of Prof. Pietro Perona. He finished his PhD in theoretical high energy physics under the supervision of Nobel laureate Prof. Gerard ‘t Hooft.
@@ Appendix A Limitations and Future Work
Limitations. One of the main limitations of this work in relation to prior work on independent component analysis and sparse coding is the unanswered question of ‘identifiability’ of latent factors in our model. Specifically, much of the early work on ICA was focused around answering to what extent the ground truth latent variables can be recovered after an unknown linear transformation [comon1994independent, 1306473]. More recently, a number of papers have demonstrated the necessary conditions for these identifiability proofs to extend to non-linear transformations of the latent variables [hyvarinen2023nonlinear]. Interestingly, these necessary conditions often revolve around temporal structure such as time-step conditioning [pmlr-v89-hyvarinen19a], or sparsity of transformations over time [klindt2021towards]. Our model, Sparse Transformation Analysis, clearly takes inspiration from these proofs by integrating these factors into a highly flexible generative modeling framework. Empirically, we demonstrate that STA does separate the ground-truth factors of variation from sequence data entirely unsupervised for a variety of datasets, including ‘real world’ datasets with more natural image statistics. We note that even in the case of models which have proved this identifiability, this is often only under strict assumptions about the true data distribution, and these assumptions are frequently seen to be invalidated on real data. These papers therefore similarly resort to measuring if their models are able to separate the underlying factors of variations in an empirical manner, as we have done in this paper. Furthermore, in our empirical analysis, we see that due to the flexibility endowed by the learned latent flows of our model, it is able to outperform models which indeed have these identifiability guarantees (such as [klindt2021towards]). That being said, in future work, we believe that it would undoubtedly be beneficial and of great interest to understand the exact necessary conditions for such a model to provably identify the generative factors. We believe such an understanding will likely be helpful for selecting optimal settings of the hyperparameters, such as the probability of switching on for the Bernoulli distribution or the scale parameter for the Laplace distribution, while similarly helping to guide the future development of unsupervised representation learning in general.
@@
Generating High-Resolution Videos. When dealing with high-resolution videos, the complexity of detailed objects and diverse backgrounds substantially increases, making it extremely challenging to effectively capture intricate motions or transformations in the highly compressed latent space. To address these challenges, incorporating additional sophisticated feature extraction methods (such as key-point tracking) could be beneficial, as demonstrated in previous motion tracking or synthesis work [sun2021multi, tulyakov2018mocogan]. Thus, the primary obstacles lie in designing powerful motion feature extraction methods and identifying meaningful latent representations that can robustly encode complex motion dynamics at higher resolutions into lower dimensions.
@@
Integration into Diffusion Models. Another critical obstacle is that standard diffusion models do not inherently define a semantically meaningful, highly compressed latent space. To integrate our methodology into diffusion-based models, we must first carefully identify an appropriate latent representation. One potential approach is leveraging the “h-space” proposed by Kwon et al. [kwon2022diffusion], which uses the bottleneck features from the score-prediction U-Net as a latent representation. However, empirical validation is still necessary to confirm whether such latent spaces are sufficiently powerful to capture transformations effectively compared to latent spaces explicitly constructed by GANs or VAEs.
@@ B.3 Proof
where ${\mathbf{V}}({\bm{z}}_{t-1})=[{\bm{v}}_{1}({\bm{z}}_{t-1}),\dots,{\bm{v}}_{K}({\bm{z}}_{t-1})]\in\mathbb{R}^{d\times K}$. From the theory of sparse dictionary learning [donoho2003optimally], if a dictionary $\mathbf{V}$} has columns that are linearly independent and the sparsity level $s$ satisfies $s<\text{spark}({\mathbf{V}})/2$, then for every vector $\delta{\bm{z}}$ that admits a $s$-sparse representation over $\mathbf{V}$}, this representation is unique. The spark of a matrix is the smallest number of columns that are linearly dependent. If $\mathbf{V}$} has full column rank and no subset of $2s$ or fewer columns is linearly dependent, then the sparse coefficients and the dictionary can be uniquely recovered (up to permutation and scaling) from sufficiently many samples.
Moreover, under the assumption A4, the dataset provides a diverse set of support patterns ${\mathbf{S}}_{t}$, ensuring that the different combinations of active vector fields are sufficiently sampled. This property is crucial for the joint recovery of the dictionary ${\mathbf{V}}({\bm{z}})$ and sparse codes ${\bm{g}}_{t}$ using methods akin to K-SVD [aharon2006k] or ER-SpUD [spielman2012exact]. When such diversity holds across $\bm{z}$}, the function-valued dictionary ${\bm{v}}_{k}({\bm{z}})$} can be recovered pointwise, as each ${\bm{z}}_{t-1}$ provides local linear constraints.

### v3

@@ 4.3 Fluid-Dynamic OT for Posterior Flow
Theorem 4.1 (Benamou-Brenier Formula (BBF) [7]).
@@ 4.4 Brownian Motion for Latent Prior Evolution
In line with [79], as we do not assume any prior knowledge of each transformation, we would like to enforce minimally informative priors. This can be achieved by considering the time evolution as Brownian motion, i.e., random trajectories. To this end, we define the potential function $\psi^{k}({\bm{z}})=-D_{k}\log p({\bm{z}}_{t})$ which advects the density $p({\bm{z}})$ through the induced velocity field $\nabla\psi^{k}({\bm{z}})$. Then according to the continuity equation, the prior evolves as:
@@ 4.5 Latent Posterior Time Evolution
TABLE II: Equivariance error $\mathcal{E}_{k}$ and average log-likelihood $\log p({\bm{x}}_{t})$ on Shapes3D [11].
@@
Figure 5: Traversals using each individual learned flow field on Shapes3D [11]. In the bracket, we indicate the transformation which the traversal results look most like. Each latent flow has separate samples per row transforming from left to right.
@@ 5.1.1 Datasets
We evaluate our method on two widely-used benchmarks for standard representation learning, namely MNIST [56] and Shapes3D [11]. For MNIST, The basic ‘pure’ transformations consist of Scaling, Rotation, and Coloring. For Shapes3D, we use the self-contained four transformation primitives, including Floor Hue, Wall Hue, Object Hue, and Scale. On both datasets, we use our spike and slab prior to generate sequences that are composed of ‘composite’ transformations.
Beyond the toy datasets, we also evaluate our method on challenging Falcol3D and Issac3D [62], two complex large-scale and real-world datasets that contain sequences of different transformations. Specifically, Falcol3D consists of indoor 3D scenes with different lighting conditions and camera positions, while Isaac3D is comprised of various robot-arm movements in dynamic environments. Since the image sequences are short, we do not consider speed variations but only enforce the spike prior to generate data sequences with sparsely-varying transformations.
We further conduct some preliminary experiments of applying our method to real-world video analysis, including autonomous driving videos on Cityscape [20] and behavior videos of social agents on CalMS [85]. Different from the used datasets above, we directly feed raw video sequences as input and let the model discover independent motions.
@@ 5.1.2 Baselines
We compare our method with some representative approaches in the field of disentangled and equivariant representation learning, including LatentFlow [79] and PoFlow [80] which adopt potential flow to evolve the latent samples, Topographic VAE (TVAE) [49] which posses topographic structured latent space, SlowVAE [52] which proposes the sparse Laplacian prior $p({\bm{z}}_{t}|{\bm{z}}_{t-1})=\prod\nicefrac{{\alpha\lambda}}{{2\Gamma(1/\alpha)}}\exp{(-\lambda|z_{t,i}-z_{t-1,i}|^{\alpha})}$, and $\beta$-VAE [38] and FactorVAE [50] which encourage the factorization of the single dimensions of latent samples. We also use the vanilla VAE [51] as a controlled baseline.
@@ 5.1.3 Metrics
We mainly evaluate the baselines using the equivariance error which is defined as $\mathcal{E}_{k}{=}\sum_{t=1}^{T}|{\bm{x}}_{t}-\texttt{Decode}({\bm{z}}_{t})|$ where ${\bm{x}}_{t}$ is the element of sequences of each transformation primitive (e.g., scaling and rotation). Since our method is unsupervised, we inspect the traversal results of each basic vector field $\nabla u^{k}+{\bm{r}}^{k}$ and select the index whose flow looks the most like the target transformation. The average log-likelihood of the sequence is also evaluated on the test set. Besides these two metrics, we also adopt the metric Variational Predictability (VP) score [106] to evaluate the disentanglement performance. Readers are kindly referred to Sec. D.3 of the supplementary for these results.
@@ 5.2.2 Quantitative Results
Table I and II present the evaluation results of the equivariance error and log-likelihood on MNIST and Shapes3D, respectively. We see STA achieves very competitive performance against other baselines. Specifically, STA outperforms all the unsupervised approaches by a large margin on equivariance error and rivals PoFlow [80] which requires supervision of each transformation primitive. Moreover, our method yields the highest log-likelihood on the test set, which is likely accounted for by the fact that our method incorporates a sophisticated transformation-centric prior over latent states which matches the statistics of the data. The sparse combination of multiple transformations thus can be seen as a kind of data augmentation. Among the transformations, the rotation has the smallest equivariance error. We expect this gain is largely due to the rotational vector field ${\bm{r}}({\bm{z}})$ introduced by the Helmholtz decomposition.
@@ 5.3.3 Periodic Transformations
Figure 10: Traversals using each individual learned flow field on Falcol3D [62]. In the bracket, we indicate the transformation which the traversal results look most like. Each latent flow has separate samples per row transforming from left to right. The bottom row displays the traversal result generated by the $6$’th latent flow field.
@@
TVAE [49] 11477.81 12568.32 11807.34 11829.33 11539.69 11736.78 11951.45
PoFlow [80] 8312.97 7956.18 8519.39 8871.62 8116.82 8534.91 8994.63
LatentFlow [79] 5798.42 6145.09 6334.87 6782.84 6312.95 6513.68 6614.27
@@ 5.5 Real-world Video Analysis
Figure 11: Traversals using each individual learned flow field on Issac3D [62]. In the bracket, we indicate the transformation which the traversal results look most like. Each latent flow has separate samples per row transforming from left to right.
@@
TVAE [49] 8441.65 8348.23 8495.31 8251.34 8291.70 8741.07 8456.78 8512.09
PoFlow [80] 6572.19 6489.35 6319.82 6188.59 6517.40 6712.06 7056.98 6343.76
LatentFlow [79] 3659.72 3993.33 4170.27 4359.78 4225.34 4019.84 5514.97 3876.01
@@ 5.5.1 Robot Arms and Indoor Scenes
Fig. 10 and 11 show the learned latent flows on Falcol3D and Issac3D [62], respectively. As can be seen from the figures, even on these challenging large-scale datasets, our method still allows for unsupervised disentanglement of complex real-world transformations. Table V and VII compares the equivariance error on the two datasets. Similar to the results on MNIST and Shapes3D, our method achieves very competitive performance against supervised ones. This demonstrates that the proposed sparsity priors also scale up to sequences of complex transformations.
Figure 12: Traversal results of learned latent flows on CalMS [85]. For each latent flow, we display two exemplary sequences, and the flow transforms the image from left to right.
@@ 5.5.2 Agent Behavioral Videos
TABLE VIII: Behavior classification results on CalMS [85].
Method MARS [73] B-Kind [85] Trajectory-LSTM [85] STA
@@
Figure 14: Traversals results of learned flow fields on downsampled segmentation masks of Cityscape [20]. Each latent flow transforms the image from left to right.
We apply our method to disentangle the complex social interactions of mice on Task $1$ of CalMS [85]. On this dataset, there exist three ground truth interactions, namely ‘investigation’, ‘attack’, and ‘mount’. We thus define $3$ latent flows and let the model learn the interactions from the raw videos. The images are of the resolution $128\times 128$.
@@ 5.5.3 Autonomous Driving Videos
Finally, we take a step further to evaluate our method on Cityscape [20], the challenging real-world autonomous driving videos. We take the sequences of segmentation masks as the training data and downsample the resolution to $64{\times}64$. Fig. 14 displays some exemplary traversals of different latent flows. On this dataset, there are no ground truth generative factors so we may have some reasonable interpretations according to the disentangled transformations: we may interpret latent flow 0 as turning left (the sidewalk region on the right side shrinks), latent flow $1$ as getting closer to the front car (the car region expands), latent flow $2$ as getting away from the front car (the car region shrinks and disappears), and latent flow $3$ as changing the right side from terrain to sidewalk. Notice that this is an initial attempt to apply our method to complex real-world video analysis. Nonetheless, this preliminary experiment demonstrates that our method could have real-world applicability for video understanding.
@@ 6 Conclusion
Acknowledgments
This research was supported in part by gifts from Cisco and OpenAI.
@@ Appendix A Limitations and Future Work
Limitations. One of the main limitations of this work in relation to prior work on independent component analysis and sparse coding is the unanswered question of ‘identifiability’ of latent factors in our model. Specifically, much of the early work on ICA was focused around answering to what extent the ground truth latent variables can be recovered after an unknown linear transformation [18, 28]. More recently, a number of papers have demonstrated the necessary conditions for these identifiability proofs to extend to non-linear transformations of the latent variables [43]. Interestingly, these necessary conditions often revolve around temporal structure such as time-step conditioning [44], or sparsity of transformations over time [52]. Our model, Sparse Transformation Analysis, clearly takes inspiration from these proofs by integrating these factors into a highly flexible generative modeling framework. Empirically, we demonstrate that STA does separate the ground-truth factors of variation from sequence data entirely unsupervised for a variety of datasets, including ‘real world’ datasets with more natural image statistics. We note that even in the case of models which have proved this identifiability, this is often only under strict assumptions about the true data distribution, and these assumptions are frequently seen to be invalidated on real data. These papers therefore similarly resort to measuring if their models are able to separate the underlying factors of variations in an empirical manner, as we have done in this paper. Furthermore, in our empirical analysis, we see that due to the flexibility endowed by the learned latent flows of our model, it is able to outperform models which indeed have these identifiability guarantees (such as [52]). That being said, in future work, we believe that it would undoubtedly be beneficial and of great interest to understand the exact necessary conditions for such a model to provably identify the generative factors. We believe such an understanding will likely be helpful for selecting optimal settings of the hyperparameters, such as the probability of switching on for the Bernoulli distribution or the scale parameter for the Laplace distribution, while similarly helping to guide the future development of unsupervised representation learning in general.
@@
Generating High-Resolution Videos. When dealing with high-resolution videos, the complexity of detailed objects and diverse backgrounds substantially increases, making it extremely challenging to effectively capture intricate motions or transformations in the highly compressed latent space. To address these challenges, incorporating additional sophisticated feature extraction methods (such as key-point tracking) could be beneficial, as demonstrated in previous motion tracking or synthesis work [85, 89]. Thus, the primary obstacles lie in designing powerful motion feature extraction methods and identifying meaningful latent representations that can robustly encode complex motion dynamics at higher resolutions into lower dimensions.
@@
Integration into Diffusion Models. Another critical obstacle is that standard diffusion models do not inherently define a semantically meaningful, highly compressed latent space. To integrate our methodology into diffusion-based models, we must first carefully identify an appropriate latent representation. One potential approach is leveraging the “h-space” proposed by Kwon et al. [55], which uses the bottleneck features from the score-prediction U-Net as a latent representation. However, empirical validation is still necessary to confirm whether such latent spaces are sufficiently powerful to capture transformations effectively compared to latent spaces explicitly constructed by GANs or VAEs.
@@ B.3 Proof
where ${\mathbf{V}}({\bm{z}}_{t-1})=[{\bm{v}}_{1}({\bm{z}}_{t-1}),\dots,{\bm{v}}_{K}({\bm{z}}_{t-1})]\in\mathbb{R}^{d\times K}$. From the theory of sparse dictionary learning [25], if a dictionary $\mathbf{V}$} has columns that are linearly independent and the sparsity level $s$ satisfies $s<\text{spark}({\mathbf{V}})/2$, then for every vector $\delta{\bm{z}}$ that admits a $s$-sparse representation over $\mathbf{V}$}, this representation is unique. The spark of a matrix is the smallest number of columns that are linearly dependent. If $\mathbf{V}$} has full column rank and no subset of $2s$ or fewer columns is linearly dependent, then the sparse coefficients and the dictionary can be uniquely recovered (up to permutation and scaling) from sufficiently many samples.
Moreover, under the assumption A4, the dataset provides a diverse set of support patterns ${\mathbf{S}}_{t}$, ensuring that the different combinations of active vector fields are sufficiently sampled. This property is crucial for the joint recovery of the dictionary ${\mathbf{V}}({\bm{z}})$ and sparse codes ${\bm{g}}_{t}$ using methods akin to K-SVD [2] or ER-SpUD [83]. When such diversity holds across $\bm{z}$}, the function-valued dictionary ${\bm{v}}_{k}({\bm{z}})$} can be recovered pointwise, as each ${\bm{z}}_{t-1}$ provides local linear constraints.